# How Many Van Goghs Does It Take to Van Gogh? Finding the Imitation Threshold

### v1

@@ Abstract
Text-to-image models are trained using large datasets collected by scraping image-text pairs from the internet. These datasets often include private, copyrighted, and licensed material. Training models on such datasets enables them to generate images with such content, which might violate copyright laws and individual privacy. This phenomenon is termed imitation – generation of images with content that has recognizable similarity to its training images. In this work we study the relationship between a concept’s frequency in the training dataset and the ability of a model to imitate it. We seek to determine the point at which a model was trained on enough instances to imitate a concept – the imitation threshold. We posit this question as a new problem: Finding the Imitation Threshold (FIT) and propose an efficient approach that estimates the imitation threshold without incurring the colossal cost of training multiple models from scratch. We experiment with two domains – human faces and art styles – for which we create four datasets, and evaluate three text-to-image models which were trained on two pretraining datasets. Our results reveal that the imitation threshold of these models is in the range of 200-600 images, depending on the domain and the model. The imitation threshold can provide an empirical basis for copyright violation claims and acts as a guiding principle for text-to-image model developers that aim to comply with copyright and privacy laws.  [^1] We release the code and data at https://github.com/vsahil/MIMETIC-2.git.
@@ 1 Introduction
The progress of multi-modal vision-language models has been phenomenal in recent years (Ramesh et al., 2021; Rombach et al., 2022; Goodfellow et al., 2022), much of which can be attributed to the availability of large-scale pretraining datasets like LAION (Schuhmann et al., 2022). These datasets consist of semi-curated image-text pairs scraped from Common Crawl, which leads to the inclusion of explicit, copyrighted, and licensed material (Cavna, 2023; Hunter, 2023; Vincent, 2023; Jiang et al., 2023; Birhane et al., 2021). Training models on such images may be problematic because text-to-image models can imitate — generate images with highly recognizable features — concepts from their training data (Somepalli et al., 2023a; Carlini et al., 2023a).  This behavior has both legal and ethical implications, such as copyright infringements as well as privacy violations of individuals whose images are present in the training data without consent. In fact, a large group of artists sued Stability AI, developers of the widely-used text-to-image models, alleging that the company’s models generated images that distinctly replicated their artistic styles (Saveri & Butterick, 2023).
@@
Figure 1: An overview of FIT, where we seek the imitation threshold – the point at which a model was exposed to enough instances of a concept that it can reliably imitate it. The figure shows four concepts (e.g., Van Gogh’s art style) that have different counts in the training data (e.g., 213K for Van Gogh). As the image count of a concept increases, the ability of the text-to-image model to imitate it increases (e.g. Piet Mondrian and Van Gogh). We propose an efficient approach, MIMETIC2, that estimates the imitation threshold without training models from scratch.
Previous work has focused on detecting when generated images imitate training images (Somepalli et al., 2023a; Carlini et al., 2023a), and its mitigation (Somepalli et al., 2023b; Shan et al., 2023). In particular, researchers found that duplicate images increase the chance of memorization and imitation. However, the relation between a concept’s prevalence in the training data and the model’s ability to imitate it remains unexplored.
@@
In this work, we ask how many instances of a concept does a text-to-image model need to be trained on to imitate it, where concept refers to a specific person or a specific art style, for example. Establishing such an imitation threshold is useful for several reasons. First, it offers an empirical basis for copyright infringement and privacy violation claims, suggesting that if a concept’s prevalence is below this threshold, such claims are less likely to be true (Saveri & Butterick, 2023; Vincent, 2023; Ceoln, 2022). Second, it acts as a guiding principle for text-to-image models developers that want to avoid such violations. Finally, it reveals an interesting connection between training data statistics and model behavior, and the ability of models to efficiently harness training data (Udandarao et al., 2024; Carlini et al., 2023b). We name this problem FIT: Finding the Imitation Threshold, and provide a schematic overview of this problem in Figure 1.
@@
The optimal methodology to measure the imitation threshold requires training multiple models with varying number of images of a concept and measuring the ability of the these models to imitate it. However, training even one of these models is extremely expensive (Bastian, 2022). We propose a tractable alternative, Measuring Imitation ThrEshold Through Instance Count and Comparison (MIMETIC2), that estimates the threshold without incurring the cost of training multiple models. We start by collecting a large set of concepts (e.g., various kinds of art styles) per domain (e.g., domain of art styles), and use a text-to-image model to generate images for each concept. Then, we compute the imitation score of each concept by comparing the generated images of that concept to its training images, and also estimate each concept’s frequency in the training data. Finally, by sorting the concepts based on increasing frequency, we estimate the imitation threshold for that domain using a change detection algorithm Killick et al. (2012).
@@
As we operate with observational data, a naive approach may be confounded by several factors, such as the quality of the imitation scoring model on different groups within the domain, or estimating the training frequencies of concepts (e.g., simple counts of ‘Van Gogh’ in the captions results in a biased estimate since the artist may be mentioned in the caption without their painting in the corresponding image). We carefully tailor MIMETIC2 to minimize the impact of such confounders (§5).
@@
Overall, we formalize a new problem – Finding Imitation Threshold (FIT; §3), and propose a method, MIMETIC2, that efficiently estimates the imitation threshold for text-to-image models (§5). We use our method to estimate the imitation threshold for two domains on four datasets, three text-to-image models that were trained on two pretraining datasets (§4). We find the imitation thresholds to range between 200 to 600 images, providing concrete insights on models’ imitation abilities (§6).
@@ 2 Background
Text-to-Image Models Generating images from textual inputs is a long studied problem, with diffusion models as the current state-of-the-art (Sohl-Dickstein et al., 2015; Dhariwal & Nichol, 2021). Diffusion models are generative models that learn to approximate the underlying data distribution (conditioned or unconditioned). Given a trained diffusion model, it is possible to sample synthetic images from the learned distribution by performing a sequence of denoising operations. Open-sourced models like Stable Diffusion (SD) Rombach et al. (2022) are trained using large datasets such as LAION (Schuhmann et al., 2022), a large image-caption dataset scraped from Common Crawl.
Dataset Issues and Privacy Violations The advancement in text-to-image capabilities, largely due to big training datasets, is accompanied by concerns about the training on explicit, copyrighted, and licensed material Birhane et al. (2021) and imitating such content when generating images (Cavna, 2023; Hunter, 2023; Vincent, 2023; Jiang et al., 2023). For example, Birhane et al. (2021) and Thiel (2023) found several explicit images in the LAION dataset and Getty Images found that LAION had millions of their copyrighted images Vincent (2023). Issues around imitation of training images has especially plagued artists, whose livelihood is threatened Saveri & Butterick (2023); Shan et al. (2023), as well as individuals whose face has been used without consent to create inappropriate content Hunter (2023); Badshah (2024). The imitation threshold would be a useful basis for copyright infringement and privacy violation claims in such cases.
Training Data Statistics and Model Behavior Pretraining datasets are a core factor for explaining model behavior (Elazar et al., 2024). Razeghi et al. (2022) found that the few-shot performance of language models is highly correlated with the frequency of instances in pretraining dataset. Udandarao et al. (2024) bolster this finding by demonstrating that the performance of multimodal models on downstream tasks is strongly correlated with a concept’s frequency in the pretraining datasets. In addition, Carlini et al. (2023b) show that language models more easily memorize highly duplicated sequences. We find a similar phenomenon: increasing the number of images of a concept increases the similarity between its generated and training images. Crucially, instead of measuring memorization, we measure imitation to find the imitation threshold. Typically, memorization refers to the replication of a specific training image whereas we use imitation to refer to the generation of a concept when a model is given different prompts, such that the concept is recognizable in the generated images.
@@ 3 Problem Formulation and Overview
Finding the Imitation Threshold (FIT) seeks to find the minimal number of images with a concept (e.g., a specific art style) a text-to-image model has to see during training in order to imitate it. FIT’s setup involves a training dataset $\mathcal{D}\triangleq\{(\mathbf{x_{i}},\mathbf{y_{i}})\mid\mathbf{x_{i}}\in\mathcal{X},\;\mathbf{y_{i}}\in\mathcal{Y}\}_{i=1}^{n}$, composed of $n$ (image, caption) pairs, where $\mathcal{X}$ and $\mathcal{Y}$ represents space of images and text captions, respectively and a model $\mathcal{M}$ that is trained on $\mathcal{D}$ to generate an image $\mathbf{x}$ given the text-caption $\mathbf{y}$. From hereafter, $\mathcal{M}(\mathbf{y})$ represents the generated image for a provided caption $\mathbf{y}$. Let $\mathbb{I}^{j}(\mathbf{x})$ be an indicator function that indicates whether a concept $Z^{j}$ is present in an image $\mathbf{x}$. Each concept $Z^{j}$ appears $c^{j}=\sum_{i}\mathbb{I}^{j}(\mathbf{x_{i}})$ times in the dataset $\mathcal{D}$, where the dataset may contain multiple concepts $\{Z^{1},Z^{2},\ldots\}$. Lastly, $\mathcal{M}^{j}_{k}$ represents a model that is trained on a dataset where $c^{j}=k$.
@@
The imitation threshold is the minimal number of training images containing some concept $Z^{j}$ from which the model $\mathcal{M}$ generates images $\mathcal{M}(\mathbf{p}^{j})$ (for different random seeds) where $Z^{j}$ is recognizable as the concept, where prompt $\mathbf{p}^{j}$ mentions the concept $Z^{j}$. [^2] Prompts $\mathbf{p}$ are usually different from the training data captions, $\mathbf{y}$. For example, if $Z^{j}$ refers to Van Gogh’s art style, then the imitation threshold is the minimal number of training images of Van Gogh’s paintings in a dataset which, when used to train a text-to-image model, can generate paintings that have Van Gogh’s art style (i.e., imitate Van Gogh’s art style).
@@
$\textit{Imitation Threshold}^{\;j}$ $\triangleq\min\left\{k\in\{1,\dots,n\}:\mathbb{I}^{j}(\mathcal{M}^{j}_{k}(\mathbf{p}^{j}))=1\right\}$ (1)
@@
Optimal Approach. Finding the imitation threshold is a causal question; the optimal manner of answering this question is the counterfactual experiment Pearl (2009). For each concept $Z^{j}$ that we want to measure the imitation threshold for, we create training datasets $\{\mathcal{D}^{j}_{1},\mathcal{D}^{j}_{2},\dots,\mathcal{D}^{j}_{k},\dots\}$ where $k$ denotes the number of images containing $Z^{j}$ that are added to a large training dataset $\mathcal{D}$ (that initially has no images containing $Z^{j}$), and then train a model $\mathcal{M}^{j}_{k}$ on each dataset $\mathcal{D}^{j}_{k}$. Once we find a model, $\mathcal{M}^{j}_{k}$ which is able to generate images that recognizably contains the concept $Z^{j}$, but $\mathcal{M}^{j}_{k-1}$ cannot, we deem $k$ as the imitation threshold for that concept. However, due to the extreme costs of training even one text-to-image model (Bastian, 2022), this approach is impractical  [^3] Naively, it will require training $\mathcal{O}(\log m)$ models for concept $Z^{j}$ (where $m$ is the total number of available images of concept $Z^{j}$), with an estimated cost of \$10M for $m=100,000$ (see Appendix M). .
@@
MIMETIC2. Instead, we propose an approach that is tractable and estimates the imitation threshold under certain assumptions (discussed later). The key idea is to use observational data instead of training multiple models with a different number of images of a concept. Such an approach has been previously used to answer causal questions, inter alia, Pearl (2009); Lesci et al. (2024); Lyu et al. (2024). Concretely, we collect several concepts from the same domain (e.g., art styles) while ensuring that these concepts have varying image counts in the training dataset $\mathcal{D}$ of a pretrained model $\mathcal{M}$. Then, we identify the count where the model $\mathcal{M}$ starts imitating concepts, and deem this count to be the imitation threshold for that domain. Note that, this threshold is domain specific and not concept specific, where a domain is defined as a set that contains several concepts belonging to the same abstract set as the specific concept for which we want to measure the imitation threshold. E.g., if concept refers to Van Gogh’s art style, then a domain would refer to a set of art styles.
To evaluate the imitation ability of a model, we build two functions, concept-count and imitation-score that counts the number of images of a concept and measures the imitation of a concept in the generated images, respectively. For a domain like art styles, we collect a set of concepts $\{Z^{1},Z^{2},\ldots\}$ (e.g., a list of various art styles), estimate each concept’s count in training data $\mathcal{D}$, and measure the imitation score for each concept using their generated and training images. Sorting the concepts based on their increasing image counts, and using a standard change detection algorithm on the imitation scores, gives us the imitation threshold for that domain. We provide the implementation details in Section 5.
@@
Assumptions. To estimate the imitation threshold using observational data, we make three assumptions. First, we assume distributional invariance between the images of all concepts in a domain. Under this assumption, measuring the imitation score of a concept $Z^{j}$ for a counterfactual model $\mathcal{M}_{k^{\prime}}$ that is trained with $k^{\prime}$ images of $Z^{j}$ is equivalent to measuring the imitation score of another concept $Z^{i}$ that currently has $k^{\prime}$ images in the already trained model $\mathcal{M}$. This assumption helps us estimate the causal effect of training a model with a different number of images of a concept, without retraining models (we empirically test the validity of this assumption in Appendix B, and we find it to hold true). Second, we assume that there are no confounders between the imitation score and the image count of a concept. These assumptions are a standard practice when answering causal questions using observational data Pearl (2009); Lesci et al. (2024); Lyu et al. (2024). Third, we assume each image of a concept contributes equally to the learning of the concept. This assumption is commonplace in sample complexity works (Valiant, 1984; Udandarao et al., 2024; Wen et al., 2024; Wang et al., 2017). These assumptions might not hold true in general, however, they are crucial for estimating the imitation threshold in a tractable manner and have been used in prior works. We also tailor our experimental setup such that these assumptions are adhered to as much as possible (§4).

### v2

# How Many Images Does It Take? Estimating Imitation Thresholds in Text-to-Image Models

@@ Abstract
Text-to-image models are trained using large datasets of image-text pairs collected from the internet. These datasets often include copyrighted and private images. Training models on such datasets enables them to generate images that might violate copyright laws and individual privacy. This phenomenon is termed imitation – generation of images with content that has recognizable similarity to its training images. In this work we estimate the point at which a model was trained on enough instances of a concept to be able to imitate it – the imitation threshold. We posit this question as a new problem and propose an efficient approach that estimates the imitation threshold without incurring the colossal cost of training these models from scratch. We experiment with two domains – human faces and art styles, and evaluate four text-to-image models that were trained on three pretraining datasets. We estimate the imitation threshold of these models to be in the range of 200-700 images, depending on the domain and the model. The imitation threshold provides an empirical basis for copyright violation claims and acts as a guiding principle for text-to-image model developers that aim to comply with copyright and privacy laws. Website: https://how-many-van-goghs-does-it-take.github.io/. Code: https://github.com/vsahil/MIMETIC-2.
@@ 1 Introduction
The progress of text-to-image models in recent years (Ramesh2021-dalle-paper; stable-diffusion-model-paper; GAN-paper-1; yao2025sv4d) is much attributed to the availability of large-scale pretraining datasets like LAION (laion-5b-paper). These datasets consist of semi-curated image-text pairs scraped from Common Crawl, which leads to the inclusion of explicit, copyrighted, and private material (washington-post-AI-artist; washington-post-AI-porn; verge-getty-sues-Stability-AI; timnit-paper-impact-on-artists; birhane2021multimodal). Training models on such images may be problematic because text-to-image models can imitate — generate images with highly recognizable features from their training data (somepalli-diffusion-model-copying; carlini2023-extracting-from-diffusion-model). This behavior has both legal and ethical implications, such as copyright infringements as well as privacy violations of individuals whose images are present in the training data without consent, that has led to several lawsuits by artists against such model providers (artists-sued-SD).
@@
Figure 1: An overview of the imitation phenomenon where we seek the imitation threshold – the point at which a model was exposed to enough instances of a concept that it can reliably imitate it. The figure shows four concepts (e.g., Van Gogh’s art style) that have different counts in the training data (e.g., 213K for Van Gogh). As the image count of a concept increases, the ability of the text-to-image model to imitate it increases (e.g., Piet Mondrian’s and Van Gogh’s art styles have higher imitation). The imitation threshold represent the number of instances a model has to be trained on such that humans recognize such concept in generated images.
Previous work proposed methods for detecting when models memorize training images (somepalli-diffusion-model-copying; carlini2023-extracting-from-diffusion-model), and mitigation techniques (somepalli2023understanding-and-mitigating; Ben-zhao2023-glaze). For instance, researchers found that duplicate images increase the chance of memorization. Typically, memorization refers to the replication of a specific training image. Instead of measuring memorization, we focus on imitation - a broader and under-explored sense of memorization, where a concept is recognizable from a generated image.
@@
In this work, we ask how many instances of a concept does a text-to-image model need to be trained on to imitate it, where concept can refer to a specific person or a specific art style, for example. Establishing such an imitation threshold is useful for several reasons. First, it offers an empirical basis for copyright and privacy violation claims, suggesting that if a concept’s prevalence in the training data is below this threshold, a model trained on such data cannot reproduce such concept (artists-sued-SD; verge-getty-sues-Stability-AI; copyrighted-image-generation). Second, it acts as a guiding principle for text-to-image models developers that want to avoid such violations. Finally, it reveals an interesting connection between training data statistics and model behavior, and the ability of models to efficiently harness training data (udandarao2024-log-trend; Carlini2022-log-trend). We posit this question as a new problem, and provide its schematic overview in Figure 1.
@@
To find the gold imitation threshold one has to train counterfactual models while varying the number of images of a concept and measuring models’ imitation abilities. However, training even one of these models is extremely expensive (orig-cost-training-t2i-model). Instead, we propose a tractable alternative, Measuring Imitation ThrEshold Through Instance Count and Comparison (MIMETIC2), that estimates the threshold without training multiple models using observational data. We start by collecting a large set of concepts (e.g., various kinds of art styles) per domain (e.g., domain of art styles), and use a text-to-image model to generate images for each concept. Then, we compute the frequency (in the training data) and the imitation score of each concept (§4.1). Then, we estimate the imitation scores using dedicated embeddings that capture the similarities of concepts (§4.2). Finally, we estimate the imitation threshold for each domain using a change detection algorithm (change-detection-paper)4.3).
@@
Operating with observational data, a naive approach may lead to a biased estimate as confounding variables such as the quality of the imitation scoring model on different groups within the domain, or estimating the training frequencies of concepts (e.g., simple counts of ‘Van Gogh’ in the captions results in a biased estimate since the artist may be mentioned in the caption without their painting in the corresponding image). We carefully tailor MIMETIC2 to deconfound such variables (§4). Since our approach is observational in nature, we also conduct counterfactual experiments to validate our results (§C). We train models on new concepts that were not seen in the training data, and vary the concept frequency to several values both smaller and larger than our estimated imitation threshold. These experiments demonstrate that our estimated imitation thresholds are accurate (with a difference of upto 10% in the estimated thresholds) .
@@
Overall, we propose and formalize a new problem that estimate the number of images of a concept that a text-to-image model requires for imitating it (§3), and propose a method MIMETIC2 that efficiently estimates the imitation threshold for these models (§4). We apply our approach to the domains of human face and art style imitation on four text-to-image models and estimate the imitation thresholds for such models to be in the range of 200-700 images (§6). These imitation thresholds provide concrete insights on imitation abilities, can act as an empirical basis for copyright violation claims, and as guiding principles for model developers.
@@ 2 Background
Text-to-Image Models Generating images from textual inputs is a long studied problem, with diffusion models as the current state-of-the-art (diffusion-model-origpaper; diffusion-model-good1).
Diffusion models are generative models that learn to approximate the underlying data distribution. Given a trained diffusion model, it is possible to sample synthetic images from the learned distribution by performing a sequence of denoising operations. Open-sourced models like Stable Diffusion (SD) (stable-diffusion-model-paper) are trained using large datasets such as LAION (laion-5b-paper), a large image-caption dataset scraped from Common Crawl.
Dataset Issues and Privacy Violations The advancement in text-to-image capabilities, largely due to large datasets, is accompanied by concerns about the training on explicit, copyrighted, and private material (birhane2021multimodal) and imitating such content when generating images (washington-post-AI-artist; washington-post-AI-porn; verge-getty-sues-Stability-AI; timnit-paper-impact-on-artists; he2024fantastic-beast-generation). For example, (birhane2021multimodal) and (Thiel2023-laion-csam) found several explicit images in the LAION dataset, and Getty Images found millions of their copyrighted images in LAION (verge-getty-sues-Stability-AI). Issues around imitation of training images has especially plagued artists, whose livelihood is threatened (artists-sued-SD; Ben-zhao2023-glaze), as well as individuals whose face has been used without consent to create inappropriate content (washington-post-AI-porn; deepfake-issue). The imitation threshold would be a useful basis for copyright infringement and privacy violation claims in such cases.
Training Data Statistics and Model Behavior Pretraining datasets are a core factor for explaining model behavior (elazar2023-WIMBD). (razeghi2022-pretraining-impact) found that the few-shot performance of language models is highly correlated with the frequency of instances in pretraining dataset. (udandarao2024-log-trend) bolster this finding by demonstrating that the performance of multimodal models on downstream tasks is strongly correlated with a concept’s frequency in the pretraining datasets. In addition, (Carlini2022-log-trend) show that language models more easily memorize highly duplicated sequences. As such, expanding the duplicates finding to the same concept re-appearing in different images, we can intuitively conjecture such concepts will be learned and then imitated by models. In this work, we study this question, and quantify the number of images to appear in the training data that are required to imitate a concept.
@@ 3 Problem Formulation and Overview
In this work, we seek to find the minimal number of images of a concept a text-to-image model requires in order to imitate it.
Defining Concept: We follow the same definition of concepts as previous work (udandarao2024-log-trend): we consider specific instances of human faces and artist styles as distinct concepts. As such, each item in a domain constitutes a concept; human individuals and artistic styles in the human face and art style domains, respectively. While there might be some visual similarities between concepts, especially for art styles, we employ discriminator models that distinguish between concepts with high accuracy (which is necessary for the accurate estimation of the imitation threshold).
@@
Setup: Our setup involves a training dataset $\mathcal{D}\triangleq\{(\mathbf{x_{i}},\mathbf{y_{i}})\mid\mathbf{x_{i}}\in\mathcal{X},\;\mathbf{y_{i}}\in\mathcal{Y}\}_{i=1}^{n}$, composed of $n$ (image, caption) pairs, where $\mathcal{X}$ and $\mathcal{Y}$ represents space of images and text captions, respectively, and a model $\mathcal{M}$ that is trained on $\mathcal{D}$ to generate an image $\mathbf{x}$ given the text-caption $\mathbf{y}$. $\mathcal{M}(\mathbf{y})$ denotes the generated image for a provided caption $\mathbf{y}$. Let $\mathbb{I}^{j}(\mathbf{x})$ be an indicator function that indicates whether a concept $Z^{j}$ is present in an image $\mathbf{x}$. Each concept $Z^{j}$ appears $c^{j}=\sum_{i}\mathbb{I}^{j}(\mathbf{x_{i}})$ times in the dataset $\mathcal{D}$, where the dataset may contain multiple concepts $\{Z^{1},Z^{2},\ldots\}$. Lastly, $\mathcal{M}^{j}_{k}$ denotes a model trained on a dataset where $c^{j}=k$.
@@
The imitation threshold is the minimal number of training images containing a concept $Z^{j}$ that model $\mathcal{M}$ generates images $\mathcal{M}(\mathbf{p}^{j})$ (with different random seeds) from a prompt $\mathbf{p}^{j}$ which mentions the concept $Z^{j}$, and $Z^{j}$ is visually recognizable.  [^1] Prompts $\mathbf{p}$ are usually different from the training data captions, $\mathbf{y}$. For example, if $Z^{j}$ refers to Van Gogh’s art style, then the imitation threshold is the minimal number of training images of Van Gogh’s paintings in a dataset used to train a text-to-image model, for which the model can generate images that imitates Van Gogh’s art style. We consider a generated image to be an imitation of a concept if the similarity between the generated image and the original images of that concept in the training data is above a threshold. We measure the similarity using a concept specific detector model and we determine the detection threshold by conducting experiments on the original images of the concept (Section˜4.2).
@@
$\textit{Imitation Threshold}^{\;j}\triangleq\min\left\{k\in\{1,\dots,n\}:\mathbb{I}^{j}(\mathcal{M}^{j}_{k}(\mathbf{p}^{j}))=1\right\}$
@@
Optimal Approach Finding the imitation threshold is a causal problem; The gold threshold can be achieved by performing the counterfactual experiment (pearl2009causality): For each concept $Z^{j}$, create $k$ training datasets $\{\mathcal{D}^{j}_{1},\mathcal{D}^{j}_{2},\dots,\mathcal{D}^{j}_{k},\dots\}$, and train a model $\mathcal{M}^{j}_{k}$ on each dataset $\mathcal{D}^{j}_{k}$. Once we find a model, $\mathcal{M}^{j}_{k}$ which is able to generate images where the concept $Z^{j}$ is recognizable, but $\mathcal{M}^{j}_{k-1}$ cannot, we deem $k$ as the imitation threshold for that concept.  [^2] In an ideal world, we can train $\mathcal{O}(log(k))$ models, with an estimated cost of \$10M if $k=100,000$ (see Appendix O). However, due to the high costs of training even one text-to-image model (orig-cost-training-t2i-model), this approach is impractical.
@@
MIMETIC2 Instead, we propose a tractable approach which efficiently estimates the imitation threshold while relying on certain assumptions (discussed at the end of this section). The key idea is to use observational data instead of training multiple models with a different number of images of a concept. This idea is a common approach to answer causal questions, when performing interventions is expensive, or unethical, inter alia, (pearl2009causality; lesci-etal-2024-causal; lyu2024causal). Concretely, we collect several concepts from the same domain (e.g., art styles) with varying image counts in the training dataset $\mathcal{D}$ of a pretrained model $\mathcal{M}$. Then, we identify the count where the model $\mathcal{M}$ starts imitating concepts, and deem this count to be the imitation threshold for that domain. Note that, this threshold is domain specific and not concept specific, where a domain is defined as the abstract set that contains the specific concept we want to measure imitation threshold for, e.g., if concept refers to Van Gogh’s art style, then its domain would be art styles. Therefore in this setup, MIMETIC2 can estimate the imitation threshold for the domain of art styles, but it cannot estimate it for Van Gogh’s art style specifically.
@@
Assumptions To estimate the imitation threshold using observational data, we make three assumptions. These are standard assumptions in the field, which are necessary for answering such complex questions. Crucially, we are able to empirically validate all of our assumptions, and find they hold for our datasets and models (Appendix˜C). First, we assume distributional invariance between the images of all concepts in a domain. Under this assumption, measuring the imitation score of a concept $Z^{j}$ for a counterfactual model $\mathcal{M}_{k^{\prime}}$ that is trained with $k^{\prime}$ images of $Z^{j}$ is equivalent to measuring the imitation score of another concept $Z^{i}$ that currently has $k^{\prime}$ images in the already trained model $\mathcal{M}$
$\vskip-5.0ptImitation\ Score(\mathcal{M}_{k^{\prime}}^{j}(Z^{j}))\approx Imitation\ Score(\mathcal{M}_{k^{\prime}}^{i}(Z^{i}))$
This assumption allows us to use observational data to estimate the imitation threshold without training models from scratch.
Second, we assume that there are no confounders between the imitation score and the image count of a concept, i.e, the imitation score of a concept is not affected by the presence of other concepts in the training data. Since there are visual similarities between concepts, it might seem that such concepts might influence the imitation score of a concept – however, our discriminator models are able to distinguish between concepts with high accuracy, alleviating the concern for such a phenomenon.
Third, we assume each image of a concept contributes equally to the learning of the concept. We further discuss these assumptions and provide empirical evidence they hold for the datasets we experiment with in Appendix˜C.
4 Proposed Methodology: MIMETIC2
We illustrate our proposed methodology in Figure˜2. At a high level, for a specific domain (e.g., human faces), MIMETIC2 estimates the frequency of each concept in the pretraining data (§4.1) and estimates the model’s ability to imitate it (§4.2). We then sort the concepts based on their frequencies, and find the imitation threshold using a change detection algorithm (§4.3).
4.1 Computing Concept Frequencies
Challenges Determining a concept’s frequency in a multimodal dataset can be achieved by employing a high-quality classifier for that concept over every image and counting the number of detected images. However, given the scale of modern datasets with billions of images, this approach is expensive and prone to classification errors.
Estimating Concept Frequency
We first make a simplifying assumption that a concept is present only if an image’s caption mentions it. We empirically test the validity of this assumption and find it to be accurate (Appendix˜D). However, this is not sufficient because concepts often do not appear in the corresponding images, even when they are mentioned in the captions. For instance, Figure˜3 showcases images whose captions contain “Mary Lee Pfeiffer”, but such images do not always include her. On average, in our experiments we find that concepts occur only in 60% of the images whose caption mentions the concept.
To address this problem, we start by retrieving all images whose caption mentions the concept of interest and then further filter out the images that do not contain the concept, as detected by a classifier (we build a classifier that indicates the presence of a particular concept in the image). We retrieve these images using WIMBD (elazar2023-WIMBD), a search tool based on a reverse index that efficiently finds documents (captions) in a dataset containing the search query (concept). To build a classifier for each concept we construct a set of high quality reference images. For example, a set of images with only the face of a single person (e.g., Brad Pitt). We collect these images automatically using a search engine (we use images from a search engine because several of the concepts in our experiments have a low number of images that cannot be used to build a classifier), followed by a manual verification to vet the images (see Appendix˜E for details). These images are used as gold reference for automatic detection of these concepts in the images. We collect up to ten reference images per concept.
Figure 3: LAION captions that mention ‘Mary Lee Pfeiffer’, the mother of Tom Cruise. She is not always present in the images (the rightmost image).
To classify whether a candidate image contains the concept of interest or not, we embed the candidate image and the concept’s reference images using an image encoder (discussed in §4.2) and measure the similarity between the embeddings. If the similarity between a candidate image and any of the reference images is above some threshold, we consider that candidate image to contain the concept of interest. This threshold is established by measuring the similarity between images of the same concepts and images of different concepts which maximizes the true positives, and minimizes false positives. We provide additional details on determining the thresholds per domain in Appendices˜G and F.
Finally, we run the classifier on all candidate images whose caption mentions the concept, and take those that are classified as positive. For each concept, we retrieve up to 100K candidate images from the pretraining dataset. We use the ratio of positive predictions to the total number of retrieved candidate images, and multiply it by the total caption count of a concept in the dataset and use that as the concept’s frequency estimate. For concepts with less than 100K candidate images, we count the number of images that are positively classified. Note that some URLs from the pretraining datasets we use are dead, a common phenomenon for URL based datasets (“link rot” (link-rot-reference1; link-rot-reference2)). On average, we successfully retrieved 74% of the candidate images.
4.2 Computing the Imitation Score
Challenges Computing the imitation score entails determining how similar a concept is in a generated image compared to its source images from the training data. Several approaches were proposed to accomplish this task, such as FID and CLIPScore (Inception-Score; Heusel2017-FID-Score; Gan-evaluation-precision-recall; hessel2022clipscore; Podell2023-SDXL-paper). To measure similarity, these approaches compute the similarity between the distributions of the embeddings of the generated and training images of a concept. The embeddings are obtained using image embedders like Inception model in case of FID and CLIP in case of CLIPScore. These image embedders often perform reasonably well in measuring similarity between images of common objects which constitutes most of their training data. However, they cannot reliably measure the similarity between two very similar concepts like the faces of two individuals or art style of two artists (somepalli2024-style-similarity; hessel2022clipscore; ahmadi2024-metric-robustness; jayasumana2024rethinkingFID). Therefore, MIMETIC2 uses domain-specific image embedders to measure similarity between two concepts from that domain.
We embed the generated and real images of concepts using these embedders and use similarity to measure a continuous imitation score. To establish a threshold, we use a change detection algorithm (described next; §4.3). At a high level, given a sequence of imitation scores, the algorithm finds the threshold on the scores above which the score is considered a (binarized) imitation of the said concept. As we are interested in finding a (binary) threshold, we binarize the score, which we find to correspond to the human perception imitation of a concept. To verify the validity of the binarized scores we conduct human subject experiments and find high correlation between human perception of imitation and the binarized threshold (Section˜6). We leave to future work the investigation of the continuous imitation score beyond the threshold.
Estimating Imitation Score We use a face embedding model (deng2018-insightface3) for measuring face similarity and an art style embedding model (somepalli2024-style-similarity) for measuring art style similarity. We find that even the specific choice of these models is crucial; for instance, in early experiments we used Facenet (facenet-paper), and observe it struggles to distinguish between individuals of certain demographics, causing drastic differences in the imitation scores between demographics. We provide more details on these early experiments in Appendix˜M, and show that our final choice of embedding models work well on different demographics.
To measure the imitation score we embed the generated images and filtered training images of a concept (obtained in §4.1) using the domain specific image embedder, and report the imitation score as the average imitation between all generated and top-10 most similar training images (cosine similarity between embeddings of generated and training images). To ensure that the automatic measure of similarity correlates with human perception, we also conduct experiments with human subjects and measure the correlation between the similarities obtained automatically and in the human subject experiments. We find this correlation to be high for both domains we experiment with (§6). We also do human subject experiments to verify the correctness of the binarized threshold and find this correlation to be high as well (§6).
4.3 Detecting the Imitation Threshold
After computing the frequencies and the imitation scores for each concept, we sort them in an ascending order of their image frequencies. This generates a sequence of points, each of which is a pair of image frequency and the imitation score of a concept. We apply a standard change detection algorithm, PELT (change-detection-paper), to find the image frequency where the imitation score significantly changes. Change detection is a classic statistical approach to find the points where the mean value of a sequence (e.g. stock value or network traffic) changes significantly. We choose PELT because of its linear time complexity in computing the change point (change-detection-algorithms-survey). We choose the first change point as the imitation threshold (see Appendix˜J for details about all change points). The application of change detection assumes that increasing the image counts beyond a certain threshold leads to a large jump in the imitation scores, and we find this assumption to be accurate in our experiments (Figure˜5(a)).
To obtain error bounds for our thresholds we perform a permutation test by sampling a subset of concepts per domain and dataset, and compute the threshold for the sampled set. We repeat this 1,000 times and report the mean and standard error of the thresholds.